### v2

# An Item is Worth a Prompt: Versatile Image Editing with Disentangled Control

Abstract
Building on the success of text-to-image diffusion models (DPMs), image editing is an important application to enable human interaction with AI-generated content. Among various editing methods, editing within the prompt space gains more attention due to its capacity and simplicity of controlling semantics. However, since diffusion models are commonly pretrained on descriptive text captions, direct editing of words in text prompts usually leads to completely different generated images, violating the requirements for image editing. On the other hand, existing editing methods usually consider introducing spatial masks to preserve the identity of unedited regions, which are usually ignored by DPMs and therefore lead to inharmonic editing results. Targeting these two challenges, in this work, we propose to disentangle the comprehensive image-prompt interaction into several item-prompt interactions, with each item linked to a special learned prompt. The resulting framework, named D-Edit, is based on pretrained diffusion models with cross-attention layers disentangled and adopts a two-step optimization to build item-prompt associations. Versatile image editing can then be applied to specific items by manipulating the corresponding prompts. We demonstrate state-of-the-art results in four types of editing operations including image-based, text-based, mask-based editing, and item removal, covering most types of editing applications, all within a single unified framework. Notably, D-Edit is the first framework that can (1) achieve item editing through mask editing and (2) combine image and text-based editing. We demonstrate the quality and versatility of the editing results for a diverse collection of images through both qualitative and quantitative evaluations. The code is available at https://github.com/asFeng/d-edit
1 Introduction
The recent development of text-to-image diffusion generative models represents a cutting-edge approach in the realm of generative models. By gradually diffusing noise into the image, these models enable sophisticated image synthesis [15, 17, 20] while preserving semantic alignment with the text prompt. One notable application is image editing, where diffusion models offer unprecedented control over various editing tasks such as inpainting [13, 1], text-guided editing [5, 14], pixel editing [12, 2], etc. Various types of image editing can be generally evaluated based on two criteria: preservation of the original image information and fidelity/consistency with the target guidance. An effective image editing process should prioritize retaining essential information from the original image while ensuring semantic alignment with the target guidance.
To improve consistency with the target guidance, some work [26, 3, 18, 25] encodes reference images by introducing additional trainable encoders to preserve identities of the reference, and adds additional controls to DPMs using methods like ControlNet [27]. However, such methods cannot incorporate the existing text prompt control flow in DPMs and therefore require large-scale pretraining which is usually costly and domain-specific. To preserve information about the original image and improve harmonization, another line of work fixes diffusion sampling trajectory (by setting random seed or using DDIM) and achieves editing by carefully tuning text prompts [11, 10], changing part of the trajectory [9], merging trajectories [8, 24], or optimizing the latent pixel space [12, 19]. This avoids additional pretraining but either relies on careful source prompt design to match the editing region or additional optimization per edit.
In this work, we propose two key techniques aimed at enhancing the aforementioned criteria: (1) Disentangled Control: To preserve information of the original image, the editing of the target item should have minimal influence on the surrounding items. The process of controlling from prompt to image should also be disentangled, ensuring that modifications to an item’s prompt do not disrupt the control flow of the remaining items. Noticing that text-to-image interactions happen in cross-attention layers for attention-based diffusion models, we propose grouped cross-attention to disentangle the prompts-to-items controlling flow. (2) Unique Item Prompt: To improve consistency with the guidance (e.g., a reference image), each item should be associated with a unique prompt that governs its generation. These prompts typically consist of special tokens or rare words. Existing works on image personalization like Dreambooth [17] and Textual Inversion [4] have extensively investigated this concept by representing the new subject with a unique prompt, subsequently employed for image generation. Compared to them, we employ independent prompts to define different items instead of the entire image. In the ideal scenario, if every item with all its details in the image can be exactly described by a unique English word, then the user can achieve all kinds of editing simply by changing the current word to the target word.
By fully leveraging the potential of prompt uniqueness and disentangled control, we introduce a versatile image editing framework for diffusion models called Disentangled-Edit (D-Edit), a unified framework that empowers most types of image editing operations at the item level, including text-based, image-based, mask-based editing, and item removal. Specifically, As illustrated in Fig. 1, starting with the target image, we initially segment it into multiple editable items (we also call background and unsegmented regions as items in the following), each of which is associated with a prompt comprising several new tokens. The associations between prompts and items are established through a two-step finetuning process, which involves optimizing the text encoder embedding matrix and the UNet model weights. Grouped cross-attention is introduced to disentangle prompt-to-item interactions, by isolating attention calculation and value updates. Various types of image editing can then be achieved by changing prompts, items, and their associations. Then, users can achieve various types of image editing by changing corresponding prompts, masks, and items, and adjusting the associations between them. This flexibility allows for a wide range of creative possibilities and precise control over the editing process. We demonstrate the versatility and performance of our framework on four image editing tasks as mentioned above, using both stable diffusion and stable Diffusion XL. We summarize our contribution as follows:
  • We propose to establish item-prompt association to achieve item editing.
  • We introduce grouped cross-attention to disentangle the controlling flow in diffusion models.
  • We propose D-Edit as a versatile framework to support various image-editing operations at the item level, including text-based, image-based, mask-based, editing, and item removal. D-Edit is the first framework that can do mask-based editing, and perform text and image-based editing at the same time.
Figure 1: The editing pipeline of using D-Edit. The user first uploads a target image which is segmented into several items. After finetuning DPMs, the user can do various types of control with inference time, including (a) replacing the model with another using a text prompt; (b) refining imperfect details caused by segmentation; (c) moving bags to the ground; (d) replacing the handbag with another one from a reference image; (e) reshaping handbag; (f) resizing the model and handbag; (g) removing background.
2 Related Works
Trajectory-Based Editing. Because natural language cannot perfectly describe a given image, a single prompt may correspond to multiple sampling trajectories with different random seeds. SDEdit [9] achieves editing by sharing the former part of the sampling process to preserve the high-level information like layout and changing the latter part for realistic reconstruction. Diffusion inversion [11, 6, 8] inverts the reverse diffusion process and forces the original and edited trajectory to share the same sampling starting point (picky to the sampling method). Interactions between the two trajectories can then be built by sharing cross/self-attention to preserve the original identity [23, 12]. Combined with P2P [5], these methods can achieve fine-grained text-based editing, but it requires an accurate captioning prompt of the original image to reverse the diffusion, and the prompt can only change by a few words. Our methods are agnostic to sampling trajectories, don’t require any prior prompt, and support more freedom to change the prompt.
Image Identity Extraction. Because image-based editing involves additional modalities for conditioning, the natural thought is to introduce additional encoders to encode the corresponding modalities. Paint-by-example [26] trains additional MLP layers following a pretrained CLIP image encoder to encode reference image information. AnyDoor [3] design an additional identity extractor to preserve the original item identity. PCDMs [18] introduced additional layers to encode the source image and target position. Because of the introduction of the additional trainable module, these models have to train on a large dataset to perform well on more images. Our method leverages the original pretrained text-encoder and UNet to encode reference images and can therefore process a wider range of images.
Image as a Word. Representing images with special tokens has been a popular choice for image personalization. Textual Inversion [4] and Dreambooth [17] represent the original subject with new tokens or rare tokens, the embedding layers or the full model are optimized with a few personalization optimization steps. We follow this line of thought in the image editing context. Instead of learning prompts from images, we learn from items and therefore can be applied when the given image contains multiple items with different subjects. The most similar works to us are SINE [28] and Imagic [7]. SINE achieves single-image editing by combining Dreambooth-trained prompt with source prompt using classifier-free guidance. Imagic optimizes the prompt embedding to be aligned with both the input image and the target text and interpolates the learned prompt to achieve editing. Compared to these methods, our framework does not require a captioning prompt in advance and can achieve more types of controls besides text-based editing.
3 Method
In this section, we discuss the details of the D-Edit framework. We first review the basics of diffusion models and text-to-image control flow. Next, we show how to establish item-prompt association through the two-step finetuning. Then, we discuss how to utilize the editability of prompts for versatile image editing operations.
3.1 Diffusion Models
Denoising diffusion probabilistic models generate high-quality images by learning to reverse the given forward Markov chain through iterative refinement. During the forward process, it works by gradually adding Gaussian noise to the original data, deriving intermediate latent as
$\begin{split}z_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}\end{split}$ (1)
with $0=\alpha_{T}<\alpha_{T-1}<..<\alpha_{0}=1$ being the noise schedule, and $\epsilon_{t}~{}\mathcal{N}(0,\mathbb{I})$. The neural network $f_{\theta}(z_{t},t)$ (like UNet) is introduced to predict the added noise $\epsilon_{t}$. The predicted noise is then used for sampling by running the reverse process which starts from the pure Gaussian noise $z_{T}$ and ends at original data $z_{0}$. Latent Diffusion Model (LDM) is the most widely adopted diffusion model for high-resolution image generation. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, LDM operates in the encoded latent space, $z_{0}=E(I)$, and maps the sampled latent representation to the original space using the paired decoder.
Text-to-Image Control. A key factor contributing to the success of LDM is its robust ability for text-to-image generation. By introducing the additional condition $y$ as the auxiliary input to $f_{\theta}(z_{t},t,y)$, LDM can generate images according to the designed user prompt. It should be noted that such prompts are usually general textual descriptions, and the final generated image is additionally controlled by the random seed in use for sampling.
Textual prompt controls the image generation through the cross-attention process. Specifically, the given text prompt $P$ containing $W$ words is first encoded by the pretrained text encoder (e.g. CLIP [16]) $g_{\phi}$ into text embedding $c=g_{\phi}(P)\in\mathbb{R}^{W\times D_{c}}$ ($W$ is the prompt length and $D_{c}$ is the embedding dimension). It is then used as input along with the image latent $z_{t}\in\mathbb{R}^{Z\times D_{z}}$ (we abuse the notation of model input and layer input) in the UNet cross-attention layer:
$\begin{split}q&=w_{q}z_{t}\in\mathbb{R}^{Z\times D}\\ k&=w_{k}c\in\mathbb{R}^{W\times D}\\ v&=w_{v}c\in\mathbb{R}^{W\times D}\\ \end{split}\qquad\begin{split}&A=\text{softmax}(qk^{T})\in\mathbb{R}^{Z\times W}\\ &O(c,z_{t})=A\cdot v,\end{split}$ (2)
where the condition $c$ is encoded into key and value vector while the image input $z_{t}$ is encoded into query vector.
3.2 Item-Prompt Association
Figure 2: Comparison of conventional full cross-attention and grouped cross-attention. Query, key, and value are shown as one-dimensional vectors. For grouped cross-attention, each item (corresponding to certain pixels/patches) only attends to the text prompt (two tokens) assigned to it.
As shown in Eq. 2, the original LDM performs text-image interaction between every token in $c$ and every pixel in $z_{t}$ through cross-attention matrix $A$. In fact, such token-pixel interactions have been shown disentangled in nature [22, 5], and the attention matrix $A\in\mathbb{R}^{Z\times W}$ is usually sparse in the sense that each column (token) only attend to several non-zero rows (pixels). For example, during image generation, the word ”bear” has higher attention scores with pixels related to the bear region compared to the remaining region.
Inspired by the natural disentanglement, we propose to segment the given image $I$ into $N$ non-overlapped items $\{I_{i}\}_{i=1}^{N}$ using segmentation model (same segmentation applied to $z^{t}$ because of emergent correspondence [21]). A set of prompts $\{P_{i}\}_{i=1}^{N}$ is adopted to replace the original text prompt $P$. Each prompt $P_{i}$ is regarded as the textual representation of the corresponding item $z^{t}_{i}$ (Details of $P_{i}$ are discussed in Sec. 3.3). As shown in Fig. 2, we force different items $I_{i}$ to be controlled by distinct prompt $P_{i}$ by masking our other items, and therefore any prompt changes in $P_{i}$ will not influence the remaining item during the cross-attention controlling flow, which is the desired property for image editing. This results in a group of disentangled cross-attentions. For each item-prompt pair ($I_{i}$, $P_{i}$), the cross-attention can be written as
$\begin{split}&q_{i}=w_{q}z^{t}_{i}\in\mathbb{R}^{Z_{i}\times D}\\ &k_{i}=w_{k}c_{i}\in\mathbb{R}^{W_{i}\times D}\\ &v_{i}=w_{v}c_{i}\in\mathbb{R}^{W_{i}\times D}\\ \end{split}\quad\begin{split}&\text{out}(\{c_{i}\},\{z^{t}_{i}\})=\Sigma_{i=1}^{N}\text{out}_{i}(c_{i},z^{t}_{i})\\ &A_{i}=\text{softmax}(q_{i}k_{i}^{T})\in\mathbb{R}^{Z_{i}\times W_{i}}\\ &\text{out}(c_{i},z^{t}_{i})=A_{i}\cdot v_{i}\end{split}$ (3)
It should be noted that such disentangled cross-attention cannot be directly used for pretrained LDMs, and therefore further finetuning is necessary to enable the model to comprehend item prompts and grouped cross-attention.
3.3 Linking Prompt to Item
We link prompts to items with two sequential steps. We first introduce the item prompt, consisting of several special tokens with randomly initialized embeddings. Then we finetune the model to build the item-prompt association.
Figure 3: Embedding layer in the text encoder. New tokens are inserted with random initialization.
Figure 4: Operations needed for different types of image editing. Each colored item has a unique prompt p.
Prompt Injection. We propose to represent each item in an image with several new tokens which are inserted into the existing vocabulary of text encoder(s). Specifically, as shown in Fig. 4, we use 2 tokens to represent each item and initialize the newly added embedding entries using Gaussian distribution with mean and standard deviation derived from the existing vocabulary. For comparisons, Dreambooth [17] represents the image using rare tokens and perfect rare tokens should have no interference with existing vocabulary, which is hard to find. Textual inversion and Imagic insert new tokens into vocabulary where the corresponding embedding is semantically initialized by given word embeddings which describe the image. This adds additional burdens of captioning the original image. We found that it is sufficient to use randomly initialized new tokens as item prompts and such randomly initialized tokens have minimal impact on the existing vocabularies.
To associate items with prompts, the inserted embedding entries are then optimized to reconstruct the corresponding image to be edited using
$\text{min}_{e}\mathbb{E}_{t,\epsilon}\left[||\epsilon-f_{\theta}(z_{t},t,g_{\Phi}(P))||^{2}\right],$ (4)
where $e\in\mathbb{R}^{NM\times D_{\text{emb}}}$ represents the embedding rows corresponding to $N$ items each with $M$ tokens.
Model Finetuning . Optimization in the first stage injects the image concept into text-encoder(s), but cannot achieve perfect reconstruction of the original item given the corresponding prompt. Therefore, in the second stage of optimization, we optimize the UNet parameters by running optimization with the same objective function as in Eq. 4. We found that updating parameters solely within cross-attention layers is adequate, as we only disentangle the forward process of these layers rather than the entire model. It should be noted that the optimizations above are running against only one image or two images (target and reference images) if image-based editing is needed.
3.4 Editing with Item-Prompt Freestyle
After the two-step optimization, the model can exactly reconstruct the original image given the set of prompts corresponding to each item, with an appropriate classifier-free guidance scale. We then achieve various disentangled image editing by changing the prompt associated with an item, the mask of an item-prompt pair, and the mapping between items and prompts. We discuss four types of image editing operations that can be achieved by varying item-prompt relationships, summarized in Fig. 4.
Text-based Item Editing.. Because the optimized model keeps the memory of the original vocabulary, replacing an item with another one described by a text prompt can be simply achieved by changing the corresponding item prompt to the target text. The learned item prompt can also be used along with other words to form a new prompt.
Image-based Item Editing. Given two images, to replace an item in the source image with another one in the reference image, all item prompts in both images should be injected into the same model. Then the item replacement can be realized by the corresponding prompt replacement Note that this setting is more general compared to classical image-based editing scenarios where the reference is an entire image instead of an item. This is particularly useful when it is hard to find a solo focus reference.
Mask-based Item Editing. Besides replacing items by changing the prompt, we can also keep the item semantics but change the appearance of the item in the image by editing the corresponding segmentation mask. The mask editing includes moving the mask position, resizing mask, refining the mask shape, or redrawing the mask. Therefore, we consider four types of mask-based editing in the experiment including changing item position, size, shape, and post-editing refinement.
Item Removal. Items in the image can be removed by deleting the corresponding item mask and item-prompt pair from the image. The deleted region will be filled by nearby region masks and corresponding item-prompt pairs.
4 Experiments
Figure 5: D-Edit enables text-guided editing in complex images. D-Edit enables the user to select any of the items from segmentation and replace it with the target text prompt.
4.1 Text-Guided Editing
Given an input image with appropriate segmentation (no captions needed), we can select any one of the items and replace the learned prompt with the target text prompt. We show such text-guided editing results in Fig. 5. Compared to null-text inversion with Prompt-to-Prompt (P2P), D-Edit can generate more realistic details and have a more natural transition between edited and unedited areas (e.g., the connection between the bike handlebar and the woman’s hand) because of disentangled control. The editing of D-Edit is more focused on the target item while the editing of inversion with P2P overflows to other regions (painting in the second example and cheese in the third) using text. Besides, unlike most text-guided methods, D-Edit does not require a caption for the original image which is extremely useful when the scene is hard to describe.
Then we show the learned item prompts can be combined with normal text words to achieve item refining, besides item replacement. As shown in Fig. 6(a), by combining the learned prompt with adjective words, we can achieve color and texture control of specific items. The preservation of the car and cake shape details after editing indicates the established association between prompts and items through finetuning, while the add-on effect shows the good quality of the learned new prompt in the vocabulary.
In Fig. 6(b), we show the results of item personalization where we generate the image using certain item prompt(s), without using the item-prompt association. This differs from Dreambooth-style personalization in that Dreambooth lacks the capability for item-level customization unless the item is cropped and focused upon, which is usually hard in an image including multiple images. Besides, it requires more images (3-5) with captions for personalization, while our method takes one image without captions. Qualitative results show that the learned item(s) can be combined with text to generate personalized images, and the text prompt can be used to personalize the background, number, and position of the item.
Figure 6: Besides text replacement, the learned prompt (denoted as [v]) can be combined with text words to achieve refinement/editing of the target item. (a) Augment an item prompt with text words while keeping other prompts unchanged for editing. (b) Generating the entire image with certain item prompt(s) augmented with text words for personalization.
4.2 Image-Guided Editing
Figure 7: Qualitative comparison of image-guided editing. D-Edit is compared with Anydoor, Paint-by-Example, and TF-ICON, on item replacement and face swapping.
For image-guided editing, the user can select one item from the reference image and use it to replace one item in the target image. We compare the editing results with baselines including Anydoors [3], Paint-by-Example [26] and TF-ICON [8] when the reference image mainly consists of a single item. As shown in Fig. 7, Paint-by-example can naturally inpaint reference items into the target scene but it falls short in keeping the identity of the reference item (the face and bird example). Anydoors can retain more relevant details from the reference image, yet it may also incorporate undesirable elements in reference, resulting in a less harmonious blend with the target image. For example, the car’s original orientation is preserved, causing it to appear out of the parking spot in the target image. Besides, it cannot preserve the face details as in the example. Compared to these methods, D-Edit is capable of seamlessly composing objects into the target while maintaining their identities.
We show more image-based editing results of D-Edit in Fig. 8 and Appendix. D-Edit can work well when the reference image contains multiple items that may be hard to separate (like the bag in hand). Additionally, D-Edit doesn’t necessitate the reference item to closely resemble its anticipated appearance in the target image, because blending through the prompt space offers smoother transitions compared to pixel-level manipulation, and the prompt-mask correspondence helps standardize the appearance of the reference item. For example, in the Ultraman example, the reference and target Ultraman can take completely different postures (kneeling v.s. standing).
Figure 8: Image-guided editing using D-Edit. Users can select any target item in the input image and replace it with another item from the same or different reference images.
Figure 9: Different types of mask-based editing: (a) moving/swapping items; (b) reshaping a (part of) an item; (c) resizing an item.
Figure 10: Post-editing refinement can be performed when obtaining imperfect results due to imperfect segmentation.
4.3 Mask-Based Editing and Item Removal
For mask-based editing, we explore four types of operations on the target items, including moving, reshaping, resizing, and refinement. As shown in Fig. 9, D-Edit can edit the shape of the target item by simply editing the corresponding mask. Because of the mask-item-prompt association, the disentangled attention can imagine and fill the new details in the edited regions according to the given item prompt, therefore leading to natural editing results. We also show the post-editing performance in Fig. 10. This can be useful when initial masks from the segmentation model do not cover the whole item, like the missing handle of the handbag and missing straps of the backpack, which will lead to imperfect (image/text/mask-guided) editing results. D-Edit can later fix these mask details and regenerate using the same random seed, and lead to refined results.
D-Edit also enables removing items by deleting the mask-item-prompt pairs. In Fig. 11, by deleting items from the scene image one by one, the resulting blank region will be re-partitioned to nearby masks and join the corresponding item-prompt pair. D-Edit will then use such new associations to imagine the blank regions and therefore lead to reasonable filling results.
Figure 11: Results of removing items one by one from the original image.
Figure 12: Qualitative comparison of textual-guided editing with and without cross-attention disentanglement
5 Conclusion
In this work, we propose D-Edit as a versatile image editing framework based on diffusion models. D-Edit segments the given image into multiple items, each of which is assigned a prompt to control its representation in the prompt space. The image-prompt cross-attention is disentangled to a group of item-prompt interactions. Each prompt is constrained to exclusively interact with the item it governs through isolated cross-attention, therefore disentangling the cross-attention control pipelines. Item-prompt associations are built up by fintuning the diffusion model which learns to reconstruct the original image using the given set of item prompts. D-Edit accomplishes image editing by allowing users to freely change or establish associations between prompts and items, and we highlight four types of image editing operations including text-based, image-based, mask-based editing, and item removal. We showcase the quality and versatility of the editing results across a diverse range of collected images through both qualitative and quantitative evaluations. Compared to baselines, D-Edit can better preserve the information of the original image while following versatile guidance.